OpenStudy (anonymous):

Given the function f(x) = −2x + 4, find x if f(x) = 20. −8 −36 44 −12

OpenStudy (anonymous):

didn't you just ask a similar question 2 secs ago?

jhonyy9 (jhonyy9):

make -2x+4=20 -2x= 16 x= 16/(-2) x = -8 hope this will help you